an electric geyser per hour of its use .it is designed to work on the mains voltage 220v.
1.does the resistance of this device remain constant during its operation/working?
2.which of the two 100W,220V night lamps has higher resistance
1
Expert's answer
2013-08-06T08:28:17-0400
1) Resistance of electric geyser are independent of voltage, time, frequency.
(Because electric geyser is only resistor that's why phase between current and voltage is zero)
2) Two lamps have P=100 w, and V=220
P=V^2/R
R=U^2/P
R=220*220/100=484 ohm (the same resistance)